OpenStudy (anonymous):

I need help on how to simplify this square root 63 I got 7square root 3.

OpenStudy (compassionate):

√63 = √(9×7) = √9 × √7 = 3√7

OpenStudy (anonymous):

so square root 98 you would get 14 square root 7?

OpenStudy (anonymous):

do you still need help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I hope this isn't to much

OpenStudy (anonymous):